Find the accumulated amount A if the principal P is invested at the interest rate of r/year for t yr. (Round your answer to the nearest cent.)
P = $1300, r = 4%, t = 3.5, compounded semiannually

Answer by rfer(16322)

A=1300(1+0.04/2)^3.5*2
A=1300(1.02)^7
A=1300(1.14869)
A=$1493.30